Problem 17-1A Comparing costs using ABC with the plantwide overhead rate LO P1, P3, A1, A2

The following data are for the two products produced by Tadros Company.

Product A Product B
Direct materials $ 15 per unit $ 24 per unit
Direct labor hours 0.3 DLH per unit 1.6 DLH per unit
Machine hours 0.1 MH per unit 1.2 MH per unit
Batches 125 batches 225 batches
Volume 10,000 units 2,000 units
Engineering modifications 12 modifications 58 modifications
Number of customers 500 customers 400 customers
Market price $ 30 per unit $ 120 per unit


The company's direct labor rate is $20 per direct labor hour (DLH). Additional information follows.

Costs Driver
Indirect manufacturing
Engineering support $ 24,500 Engineering modifications
Electricity 34,000 Machine hours
Setup costs 52,500 Batches
Nonmanufacturing
Customer service 81,000 Number of customers


Required:

(Loss amounts should be indicated with minus sign. Round your per unit cost answers to 2 decimal places.)

Which method of product costing gives better information to managers of this company?
Answer #1

a) Estimated total mfg. overhead cost Estimated direct labor hours Predetermined overhead rate/DLH $192,000 [$24500 + $34000Volume for engineering modifications = 12 + 58 = 70; batches = 125 + 225 = 350. Volume for machine hours = (10000 units x 0.158 Activity Activity Overhead costs assigned to Product B: Rate Usage Engineering support $350 Electricity $10 2,400 Setup co